What is the given point-slope form in slope-intercept form for y-2=3(x+1)?
Answer by srinivas.g(540)

slope intercept form : y= mx+c
given equation
y-2=3(x+1)
y-2 =3*x+3*1
y-2= 3x+3
move -2 to the right
y= 3x+3+2
y= 3x+5
result : slope -intercept form y= 3x+5